How strong is the electric field between the plates of a 0.60-μF air-gap capacitor if they are 1.0 mm apart and each has a charge of 71 μC ?
Answer #1

according to capacitance formula

C= Q/ V

where V is voltage between plates

=> V= Q/C = (71*10^-6)/(0.60*10^-6)=118.33 v

electric field between plates E = V/d

where d is distance between plates , d= 1.0 mm = 10^-3 m

=> E = (118.33)/10^-3 = 118.33*10^3 v/m